A particle having mass 0.5kg is projected under gravity with a speed of 98m/s at an angle of 60°.
The magnitude of change in momentum( in N-sec) of a particle after 10 seconds is?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
 
Concider the horizontal motion of that object. There is no force is acting and momentum is conserved (I.e. change in momentum = 0).
On the vertical motion, the only force acting on the body is mg. Hence, according to second law of motion,
  Change in momentum = F*t
Change in momentum in verticle = 49 kgm/s.
Resultant2 = vx+ vy2
                    = 0 + 492
Resultant = 49 kgm/s